### v1

# Ab initio correlated calculations without finite basis-set error: Numerically precise all-electron RPA correlation energies for diatomic molecules

Abstract
In wavefunction-based ab-initio quantum mechanical calculations, achieving absolute convergence with respect to the one-electron basis set is a long-standing challenge. In this work, using the random phase approximation (RPA) electron correlation energy as an example, we show how to compute the basis-error-free RPA correlation energy for diatomic molecules by iteratively solving the Sternheimer equations for first-order wave functions in the prolate spheroidal coordinate system. Our approach provides RPA correlation energies across the periodic table to any desired precision; in practice, the convergence of the absolute RPA energies to the meV-level accuracy can be readily attained. Our method thus provides unprecedented reference numbers that can be used to assess the reliability of the commonly used computational procedures in quantum chemistry, such as the counterpoise correction to the basis set superposition errors, the frozen-core approximation, and the complete-basis-set extrapolation scheme. Such reference results can also be used to guide the development of more efficient correlation-consistent basis sets. The numerical techniques developed in the present work also have direct implications for the development of basis error-free schemes for the GW method or the ab initio quantum chemistry methods such as MP2.
Introduction.- The ab initio solution of the many-electron Schrödinger equation requires convergence with respect to both the correlated methods and the one-electron basis set. The first aspect concerns how much electron correlation is captured in a given Hilbert space, while the second aspect characterizes the completeness of the Hilbert space itself. Various correlated methods have been developed in quantum chemistry and condensed matter physics to describe the interacting many-electron systems [1, 2, 3, 4], and considerable effort has been devoted to benchmark their performance in prototypical problems [5, 6]. However, these benchmarks are usually performed within a small Hilbert space defined typically in terms of a finite Gaussian-type orbital (GTO) basis set. Convergence with respect to the one-electron basis sets is usually attained by fitting the results obtained at finite basis sets in terms of an empirical formula, attempting to reach the so-called complete basis set (CBS) limit via extrapolation. As such, the accuracy of the extrapolated CBS results necessarily depends on the availability and quality of systematically improvable one-electron basis sets and the reliability of the extrapolation rule itself. It is therefore highly desirable to find alternative ways to establish the converged limit of correlated methods without relying on the extrapolation approach.
Among various electronic structure methods, the random phase approximation (RPA) stands out as a cornerstone that bridges quantum chemistry methods and density functional theory. Practical applications show that RPA-based methods are suitable for describing delicate ground-state energy differences for both molecular and extended systems. However, a major hurdle in the RPA calculations is its very slow convergence with respect to one-electron basis set, rendering the numerically fully converged results different to obtain. Calculating numerically converged RPA atomization energies via extrapolation is highly non-trivial [7]. Here, we develop an approach that allows one to obtain numerically precise RPA electron correlation energy without suffering from the basis-set incompleteness error (BSIE). The essential idea behind this approach is to compute the non-interacting Kohn-Sham (KS) density response function, the central quantity of RPA, directly from the first-order KS wavefunctions (WFs), which themselves are determined by solving the Sternheimer equation. The key point here is that, instead of expanding the Sternheimer equation in terms of a finite basis set, we solve it on a real-space grid, arriving at a solution that can be made arbitrary accurate. With a numerically fully converged KS density-response function, one can then obtain the absolute RPA correlation energy without BSIE. This technique has previously been applied to isolated atoms [8], whereby the 3-dimensional (3D) Sternheimer equation reduces to a 1-dimensional (1D) radial differential equation, which can be conveniently solved on a dense logarithmic grid. In this work, we extend this technique to diatomic molecules by solving the 2-dimensional (2D) radial Sternheimer equation in the prolate spheroidal coordinates system. This extension enables us to compute the fully converged absolute RPA correlation energy for diatomic molecules. This extension is a critical step towards a general approach to address the BSIE of RPA and related methods. By treating diatomic molecules in a numerically precise way, we can obtain fully converged all-electron (AE) RPA binding energy curves for any diatomic molecules across the periodic table. This permits a rigorous assessment of the BSIE of the commonly used GTO basis sets in quantum chemistry, and further quantify the errors arising from the commonly used frozen-core (FC) approximation and the basis set superposition errors (BSSEs). Moreover, such reference data can also guide us in developing more efficient atomic basis sets suitable for correlated calculations, especially for extended systems where currently available GTO basis sets have severe deficiencies. Last but not least, this approach can be directly extended to scaled-opposite-spin second-order Møller-Plesset perturbation theory (SOS-MP2) and the $GW$ methods, and thus has broad implications for correlated calculations.
The Sternheimer equation in prolate spheroidal coordinates system.—The key to basis-error-free RPA calculations is to obtain numerically accurate first-order KS wavefunction (WF) induced by arbitrary external perturbation. To introduce our approach, we start with the KS Hamilitonian $H^{(0)}$,
$H^{(0)}=-\frac{1}{2}\nabla^{2}+v_{eff}(\bm{r})$ (1)
where $v_{eff}(\bm{r})$ is the KS effective potential. Upon adding a small frequency-dependent perturbation $V^{(1)}(\bm{r})e^{i\omega t}$ to the Hamiltonian $H^{(0)}$, the linear response of the system is governed by the following frequency-dependent Sternheimer equation [8],
$(H^{(0)}-\epsilon_{i}+i\omega)\psi_{i}^{(1)}(\bm{r},i\omega)=(\epsilon_{i}^{(1)}-V^{(1)})\psi_{i}(\bm{r})$ (2)
where $\psi_{i}(\bm{r})$ and $\epsilon_{i}$ are the KS orbitals and orbital energies, and $\psi_{i}^{(1)}(\bm{r},i\omega)$ and $\epsilon_{i}^{(1)}$ are their first-order variations, respectively. This 3D differential equation simplifies for diatomic molecules which have the rotational symmetry with respect to the axis connecting the two nuclei. To exploit this symmetry, we chose the prolate spheroidal coordinates system, whereby the two atoms are placed at the two poles of the ellipsoid, and the position of a point in space is described a set of variables $(\mu,\nu,\theta)$. These prolate spheroidal variables are related to the Cartesian variables as $x=R/2\text{sinh}(\mu)\text{sin}(\nu)\text{cos}(\theta)$, $y=R/2\text{sinh}(\mu)\text{sin}(\nu)\text{sin}(\theta)$, and $z=R/2\text{cosh}(\mu)\text{cos}(\nu)$, with $R$ being the distance between the two atoms, and $0\leq\mu\leq\infty$, $0\leq\nu\leq\pi$, $0\leq\theta\leq 2\pi$. A graphical illustration of the geometrical meaning of these variables and a discussion of the representation of the WFs and operators in this coordinate system are provided in the Supplemental Material (SM) (see Fig. S1 and Sec. I A). In the prolate spheroidal coordinates system, the Hamiltonian and WFs adopt the following forms,
$H^{(0)}=-\frac{1}{2}\nabla^{2}+v_{eff}(\mu,\nu)$ (3)
$\psi_{i}(\bm{r})=f_{i}(\mu,\nu)e^{im\theta}$ (4)
The simplification comes from the fact that the dependence of the physical quantities on the angular variable $\theta$ can be eliminated.
A commonly used approach to evaluate the RPA correlation energy is the resolution of identity (RI) approximation [9, 10], under which the KS response function is represented as a matrix within a set of auxiliary basis function (ABF) $\{P(\bm{r})\}$. By taking the ABFs $\{P(\bm{r})\}$ as the external perturbations, one can formulate a RI-RPA scheme that is free of single-particle BSIE, as discussed in Ref. [8]. The idea follows from a similar BSIE correction scheme for $GW$ developed in the linearized augmented plane wave framework [11, 12, 13]. The mathematical formulation of Sternheimer-based RI-RPA scheme for diatomic molecules is discussed in Sec. I B of the SM.
As shown in Ref. [8], the error due to the incompleteness of the ABFs is much smaller than that of the single-particle basis. In particular, the standard ABFs constructed on the fly in the FHI-aims code [10, 14] is sufficiently good for most practical RPA calculations. Nevertheless, this error is still visible and, to achieve a numerical precision of the absolute RPA correlation energy to the meV level, it is desirable to eliminate the BSIE of the ABFs as well. To this end, we use the iterative diagonalization method to directly determine the eigenspectra of the operator $\chi^{0}(i\omega)v$. This approach has been used to calculate the dielectric function in the pseudopotential plane-wave context [15, 16, 17]. The key difference here is that we solve the Sternheimer equation on a non-uniform grid and fully exploit the symmetry of diatomic systems. Only by doing so are converged all-electron RPA calculations possible for molecules.
Specifically, we start with a trial eigenfunction $\phi(\bm{r})$ which is set in the following form to account for the prolate spheroidal symmetry,
$\phi(\bm{r})=\phi(\mu,\nu)e^{iM\theta}\,.$ (5)
Now we need to determine the first-order density $\Delta n(\bm{r})$ upon applying $\chi^{0}v$ to the trial function $\phi(\mu,\nu)e^{iM\theta}$, i.e.,
$\chi^{0}(i\omega)v\phi(\mu,\nu)e^{iM\theta}=\Delta n(\bm{r},i\omega)=\Delta n(\mu,\nu,i\omega)e^{iM\theta}\,,$ (6)
and then take the resultant $\Delta n(\bm{r},i\omega)$ as the input vector of the next iteration. The process is repeated until $\Delta n(\bm{r},i\omega)$ is converged within a given threshold. By doing so a pair of eigenvalue and eigenvector of the $\chi^{0}v$ operator is obtained.
We further note that the action of the combined $\chi^{0}v$ operator on a function of the form given in Eq. 5 can be executed in two successive steps. First, applying the Coulomb operator $v$ on $\phi(\bm{r})$ amounts to computing the Hartree potential corresponding to a density distribution of $\phi(\bm{r})$,
$\phi_{h}(\bf{r})=\int\frac{\phi(\bf{r^{\prime}})}{|\bf{r}-\bf{r^{\prime}}|}d\bf{r^{\prime}}\,.$ (7)
In practice, $\phi_{h}(\bf{r})$ is computed by solving the Poisson equation on the prolate spheroidal coordinates system.
In the second step, consider the resultant $\phi_{h}(\mu,\nu)$ as a perturbation and solve the Sternheimer equation to determine the first-order WF $f_{i}^{(1)}(\mu,\nu,i\omega)$. Then, the variation of the electron density is given by
$\Delta n(\mu,\nu,i\omega)=\sum_{i}^{occ}f_{i}^{*}(\mu,\nu)*f_{i}^{(1)}(\mu,\nu,i\omega)+c.c\,,$ (8)
where $c.c$ denotes the complex conjugate of the former term. Following these two steps, the result by applying the $\chi^{0}v$ operator on an input vector can be obtained. $\Delta n(\mu,\nu,i\omega)$ provides the input vector for the next iteration. This process is repeated until the output and input vectors are parallel to each other. More rigorous derivations and implementation details are given in Sec. I C of the SM.
In this work, we use the package ARPACK [18] based on the Arnoldi algorithm [19], together with the above-described two-step procedure, to iteratively diagonalize $\chi^{0}(i\omega)v$. In particular, ARPACK can provide a specified number of eigenvalues arranged in a desired ascending or descending order. In the present case, $\chi^{0}(i\omega)v$ is a negative definite operator, and all eigenvalues are negative. The contribution of each eigenstate to the RPA correlation energy is positively correlated with the magnitude of the eigenvalue. Concretely, the RPA correlation energy for a diatomic molecule is given by
$E_{c}^{\mathrm{RPA}}=\frac{1}{2\pi}\int_{0}^{\infty}{{\rm d}\omega}\sum_{M=0}^{M_{max}}\sum_{i}^{N_{eigen}}[ln(1-e_{i,M})+e_{i,M}]\,$ (9)
where $e_{i,M}$ is the $i$-th eigenvalue of $\chi^{0}v$ in the $M$-th angular momentum channel. Since $M$ is a good quantum number, the eigenvectors of $\chi^{0}(i\omega)v$ can be grouped according to different $M$’s. In practical calculations, we set the parameter for ARPACK such that it will output the eigenvalues in order of the magnitude , from large to small, so that the contribution to the RPA correlation energy will become gradually smaller. Systematically increasing the number of eigenvalues in a given magnetic quantum channel $M$, one can achieve arbitrary precision for a given $M$. Then, keeping increasing the magnetic quantum number $M$ and repeating the above process, one can finally obtain the RPA correlation energy to any desired precision.
Our approach is implemented in the all-electron FHI-aims code package [20]. The employment of numerical atomic orbital (NAO) basis set enables an accurate description of the manifold of occupied states, including the core one. This is also an essential factor that makes the high-precision all-electron RPA calculations possible. Further implementation details are given in Sec. II of the SM.
Result.-We first show that the all-electron RPA correlation energy for diatomic molecules can be converged to arbitrary precision within our approach. According to Eq. 9, the maximum magnetic quantum number $M_{max}$, and the number of eigenvalues $N_{eigen}$ for each $M$ are obviously the two key parameters controlling the precision of the calculation. In addition, the number of frequency points and the real-space grid points in the prolate spheroidal coordinates system are two other parameters that affect the numerical precision. The frequency integration is the easiest and the discretization error can be easily reduced to the $\mu$eV level using the modified Gauss-Legendre grid [10] or the minimax grid [21]. This has been previously shown in the literature [22] and will not be discussed here.
Next, we take the N2 molecule (with an equilibrium bond length of $1.098$) as a concrete example to check convergence behavior of the absolute RPA correlation energy with respect to the grid size $(N_{\mu},N_{\nu})$, as well as $N_{eigen}$ and $M_{max}$. Regarding the real-space grid $(N_{\mu},N_{\nu})$, as is demonstrated in Table S1 of the SM, one can achieve a numerical precision below 1 meV using a grid size of $(90,90)$ and this can be further refined to 0.1 meV when doubling the grid size. The convergence tests for heavier elements are also provided in the SM (Tables S2 and S3). We find that numerical precision better than 1 meV can be achieved for all-electron RPA correlation energies for a real-space grid size of the order of $O(100)\times O(100)$. In subsequent calculations, we ensure that the grids are dense enough so that the absolute energy is converged within 1 meV.
Now we proceed to examine the convergence behavior with respect to the two key parameters: $N_{eigen}$ and $M_{max}$. In Table LABEL:tab:1 we present the RPA correlation energies for both the N2 molecule and the isolated N atom for increasing number of eigenvalues $N_{eigen}$, whereby $M_{max}$ is fixed at 9. For a balanced treatment, in atomic calculations only $N_{eigen}/2$ eigenvalues are included. The obtained binding energies of the N2 molecule for different $N_{eigen}$ are also presented in Table LABEL:tab:1. We see that when $N_{eigen}$ is increased from 900 to 1000, the absolute RPA correlation energies of N2 and N changes by 0.15 and 0.08 meV, respectively; the binding energy is only changed by 0.001 meV. We note that such a convergence pattern does not change with $M_{max}$.
Table 1: RPA@PBE correlation energies (in eV) for N2 and the N atom for different numbers of eigenvalues $N_{eigen}$ with $M_{max}=9$. Note that only half of $N_{eigen}$ eigenvalues is used for the N atom calculation. Here the grid density is (150,120). The binding energies $\Delta E^{\text{RPA}}_{c}=E^{\text{RPA}}_{c}(\text{N}_{2})-2E^{\text{RPA}}_{c}(\text{N})$) are presented in the fourth column.
$N_{eigen}$ $E^{\text{RPA}}_{c}(\text{N}_{2})$ $E^{\text{RPA}}_{c}(\text{N})$ $\Delta E^{\text{RPA}}_{c}$
100 -23.36237 -9.22194 -4.91849
200 -23.38841 -9.23521 -4.91800
300 -23.39446 -9.23829 -4.91789
400 -23.39687 -9.23951 -4.91785
500 -23.39808 -9.24013 -4.91782
600 -23.39879 -9.24049 -4.91781
700 -23.39923 -9.24071 -4.91780
800 -23.39954 -9.24087 -4.91780
900 -23.39975 -9.24098 -4.91779
1000 -23.39990 -9.24106 -4.91778
Table 2: Contributions to the RPA@PBE correlation energies of $N_{2}$ and $N$ from different $M$ channels. For each $M$, $N_{eigen}=1000$ and 500 are used for the $N_{2}$ and $N$ calculations, respectively. The fourth column represents the $M$-resolved contributions to the RPA correlation part of the binding energy.
$M$ $E^{\text{RPA}}_{c}(\text{N}_{2})$ $E^{\text{RPA}}_{c}(\text{N})$ $\Delta E^{\text{RPA}}_{c}$
0 -12.32344 -4.32484 -3.67376
1 -8.40979 -3.83432 -0.74116
2 -1.84721 -0.71361 -0.42000
3 -0.51048 -0.22959 -0.05129
4 -0.16809 -0.07550 -0.01709
5 -0.07027 -0.03155 -0.00717
6 -0.03426 -0.01537 -0.00351
7 -0.01860 -0.00834 -0.00192
8 -0.01093 -0.00489 -0.00114
9 -0.00682 -0.00305 -0.00072
Total -23.39990 -9.24106 -4.91778
Finally, we examine the convergence behavior of the RPA correlation energy with respect to the maximum magnetic quantum number $M_{max}$. As indicated in Eq. 9, the full RPA correlation energy can be decomposed into contributions from different $M$ channels. In Table LABEL:tab:2, we present the $M$-resolved contributions to the absolute RPA correlation energies of both N2 and N, as well as to their differences (the binding energies), up to $M=9$. It should be mentioned that, except for $M=0$, the value for each $M$ in Table LABEL:tab:2 represents the sum of the results of two degenerate channels ±$\pm M$. From Table LABEL:tab:2, it can be seen that, despite the rapid decrease in the contribution to $E_{c}^{\text{RPA}}$ as $M$ increases, a sizable contribution of 7 meV is still visible for $M=9$. Obviously, to achieve an absolute convergence of the RPA correlation energy to the meV accuracy, a much larger $M_{max}$ is needed.
In this connection, we conducted a theoretical analysis of the convergence behavior of the RPA correlation energy of diatomic molecules with respect to $M_{max}$, and found that the correlation energy should converge as $1/M_{max}^{3}$, i.e.,
$E_{c}^{\text{RPA}}(M_{max})=E_{c}^{\text{RPA}}(\infty)+\frac{\alpha}{M_{max}^{3}}\,.$ (10)
This is similar to the atomic case where the correlation energy converges as $1/L_{max}^{3}$ with $L_{max}$ being the maximum azimuthal quantum number [23, 24, 8]. An in-depth theoretical derivation of this asymptotic behavior and numerical verification are provided in Sec. III B of the SM. Figs. S2 and S3 in the SM show that the RPA correlation energies of both the N atom and N2 as a function $M_{max}$ follow perfectly the behavior given by Eq. 10 with a conference level (the $R^{2}$ value) higher than $99.995\%$. Fitting the data from $M_{max}=10$ to 16, we obtain $E_{c}^{\text{RPA}}(\infty)=-23.41679$ eV for the N2 molecule, with the coefficient $\alpha=12.45267$ eV. Thus, the residual BSIE due to contributions beyond $M_{max}$ is $12.45267/M_{max}^{3}$ eV, amounting to about 17 meV for $M_{max}=9$. To have the absolute energy converged to 1 meV, a value of $M_{max}=23$ is expected. A similar fitting for the N-atom results yields $E_{c}^{\text{RPA}}(\infty)=-9.24857$, and $\alpha=5.52887$. Now, since the absolute RPA correlation energies for both N2 and N follow the same asymptotic behavior, it immediately follows that the binding energies yielded their differences, also follow the same behavior as Eq. 10, although with a much smaller coefficient. A fitting for the binding energies yields $\Delta E_{c}^{\text{RPA}}(\infty)=-4.91965$, and a coefficient $\alpha=1.39493$ eV. Therefore, at $M_{max}=9$, the BSIE for the RPA binding energies is about 1.9 meV.
Thus, we have completely resolved the convergence problem with respect to $M_{max}$ for diatomic molecules. In practical calculations, choosing $M_{max}=9$ can yield binding energies that reach meV-level accuracy. Of course, the BSIE with respect to $M_{max}$ can be completely eliminated either by running calculations with a large $M_{max}$ (say $M_{max}=23$) or extrapolating the results to $M_{max}\rightarrow\infty$ using Eq. 10.
[…]
In Fig. S4 of the SM, we compare the BSSEs present in the traditional sum of states scheme and the present approach for three series of AO basis sets, again using N2 as an illustration example. As is evident, in the traditional approach, the BSSEs are huge and do not vanish even with the largest available basis set in each series. Such BSSEs mainly come from the imbalanced description of the unoccupied manifold of the Hilbert space of the molecule and the atom. In contrast, in the present approach, the BSSEs are vanishingly small, as expected. Apparently, in our approach, there is no need to invoke the counterpoise procedure to correct BSSEs any more.
The core-core and core-valence electron correlations usually contribute little to the binding energies, and it is a common practice to employ the FC approximation in correlated calculations. However, the actual error due to the FC approximation is often hard to rigorously assess, especially for heavy elements. This is because within conventional approaches, the commonly used atomic orbital basis sets are not completely balanced in describing the valence electrons and core electrons, the BSIE and FC error are intervened. Our approach here eliminates the BSIEs on equal footings for both AE and FC calculations and thus provides an unambiguous way to assess the FC errors. In Figs. S5 of the SM, we present the AE and FC binding energy curves for N2, P2, and As2. For P2 and As2, we also freeze different core shells and observe how the results will change. The results indicate that freezing all core electrons leads to an increase in errors from N to As (48 meV for N2, 52 meV for P2, and 160 meV for As2). However, if one just unfreezes the outermost core electrons ($1s$ for N, $2s2p$ for P, and $3s3p3d$ for As), the obtained FC results become almost indistinguishable from the AE results. This investigation clarifies how to perform proper FC calculations without compromising accuracy.
[…]
Figure 1: RPA@PBE binding energy curves of Kr2 obtained using different basis sets, in comparison with the reference curves obtained using the approach developed in this work. BSSEs are corrected for cc-pV5Z and aug-cc-pV5Z basis sets; for NAO-VCC-5Z, both CP-corrected and uncorrected results are presented. The FC approximation is used for all calculations.
Finally, we present in Table LABEL:tab:mol_dimer the accurate binding energies of a set of diatomic molecules as predicted by the Sternheimer approach developed in this work. The set of molecules are chosen to be representative, composed of elements from light to heavy and bound together via different interaction types. For light elements up to Ar, AE calculations are performed, while for elements starting with K, inner-core elements are frozen. We showed above that binding energies obtained from such FC calculations are essentially identical to AE calculations. For comparison, we also include the accurate results reported by Humer it al. in Ref. 7, obtained using the explicitly correlated dRPA-F12 method [28] with the GTO basis set. Unfortunately, such accurate results are only available for a few molecules containing light elements. Table LABEL:tab:mol_dimer shows that, with an MAE of only 0.07 kcal/mol, dRPA-F12 method and the Sternheimer approach agree with each other remarkably well, confirming the excellent performance of the former in mitigating the one-electron BSIE.
Table 3: RPA@PBE binding energy (in kcal/mol) for a set of diatomic molecules. The third column represents the results obtained using the Sternheimer approach developed in the present work. The fourth column lists very accurate results taken from the work of Humer et al [7], obtained using explicitly correlated dRPA-F12 method. The mean absolute error (MAE) is measured for molecules where the dRPA-F12 results are available.
Molecule Bond length This work Ref. [7]
H2 0.74 108.72 108.69(-0.03)
He2 4.30  0.001 /
Li2 2.70  18.84  18.91(0.07)
N2 1.10  223.40 224.48(0.08)
F2 1.43  30.40  30.61(0.21)
LiH 1.60  54.66 54.48(-0.16)
HF 0.92 132.77 132.78(0.01)
LiF 1.58 127.36 127.36(0.00)
CO 1.14 245.61 245.65(0.04)
O2 1.21 113.79 113.74(-0.05)
Ne2 3.24   0.041 /
Na2 3.18  14.23 /
P2 1.91 117.19 /
Cl2 2.02  50.18 /
NaCl 2.40  90.24 /
Ar2 3.84  0.204 /
K2 4.14  10.32 /
Cu2 2.21  39.78 /
ZnO 1.72  34.45 /
Kr2 4.10  0.303 /
As2 2.06  91.10 /
Rb2 4.50   9.18 /
Ag2 2.53  34.43 /
Au2 2.53   64.11 /
MAE / / 0.07
In Table S4 of the SM, we further present the FC-RPA binding energies for a larger set of diatomic molecules, obtained using the Sternheimer approach and conventional schemes with the largest available GTO cc-pV6Z/cc-pV5Z basis set, one still has an MAE of 1.60kcal/mol. Fortunately, by extrapolating to the CBS limit, the BSIE can be significantly reduced to below 0.4 kcal/mol, for both GTO and NAO basis sets. However, for heavy elements, such correlation-consistent basis sets are unavailable, making conventional approaches unusable to obtain high-quality RPA results. In contrast, our basis-error-free RPA results provide invaluable references for developing high-quality AO basis sets for heavy elements. This possibility will be explored in future work.
We note that the numerical technique developed in the present work for correlated calculations without basis errors does not only apply to RPA, but is also directly extendable to any method based on density response functions $\chi^{0}$. One prominent example is the SOS-MP2 approach [29, 30], which can be calculated as
$E_{\text{SOS-MP2}}=-C_{\text{OS}}\int_{0}^{\infty}\frac{\mathrm{d}\omega}{2\pi}\operatorname{Tr}\left[v\chi_{0}^{\uparrow}v\chi_{0}^{\downarrow}\right]$ (11)
where $\chi_{0}^{\sigma}$ with $\sigma=\uparrow,\downarrow$ is spin-resolved density response function and $C_{OS}=1.3$. The SOS-MP2 has been found to produce better results compared to the standard MP2 in certain situations [30] without the need of calculating the second-order exchange component. In Sec. VII of the SM, we show that numerically highly-precise SOS-MP2 correlation energy can be obtained using our technique, similar to the RPA case.
Summary.-In this work we developed a numerical approach that allows to perform basis-error-free calculations ab initio correlated methods that are formulated in terms of density response function. The present implementation is restricted to the RPA and SOS-MP2 methods and diatomic molecules, but extending this approach to more general situations is foreseeable. The approach yields numerically precise absolute all-electron correlation energies and applies to both light and heavy elements. It thus offers an unambiguous way to quantify the numerical uncertainty associated with the commonly used computational protocol in quantum chemistry, i.e., finite GTO basis set with extrapolation to the CBS limit, frozen-core approximations, and the counterpoise correction to the BSSE. The reference results can also be used to guide the development of more efficient correlation-consistent AO basis sets, particularly for cases where such basis sets are not yet available.